The acetic acid solution is 1.49 x 10 -3 ± 4 x 10 -5 M. Report the pH of this solution and its uncertainty
Answer #1

As dissociation of acetic acid will give ,CH3COO- and H+

Hence its pH = - log (H+)  

pH = - log (1.49 x 10 -3) = - log(1.49) - log 10 -3 = -0.1731 + 3 = 2.82

*For y = log (a ±S(a)), uncertainty in y i.e .S(y) = 0.434*(S(a)/a)

Hence,

uncertainty in pH will be 0.434 *(1.49 x 10 -3 /4 x 10 -5) = .01165

Hence,

pH of solution with uncertainty = 2.82 ± 0.01165